OpenStudy (anonymous):

HELP ASAP I GIVE MEDALS The surface area of a cube is 486 ft2. What is the area of one face of this cube

OpenStudy (anonymous):

Well, the formula for surface area of a cube is 6a squared where a is the length of one side of the cube.

OpenStudy (anonymous):

Is the answer 18.

OpenStudy (anonymous):

So in order to find the area of one side you would want to set 6a^2=6 and solve for a^2.

OpenStudy (anonymous):

Not equal to 6 sorry.

OpenStudy (anonymous):

Equal to 486

OpenStudy (anonymous):

No, 18 is not the right answer.

OpenStudy (anonymous):

what is the answer plz tell me.

OpenStudy (anonymous):

You need to solve the equation 6a^2=486 for a^2. Thats your answer.

OpenStudy (anonymous):

i dont know how can you tell me the answer plz

OpenStudy (anonymous):

You need to isolate the a^2. wat can you divide by both sides to get the a^2 alone?

OpenStudy (anonymous):

I dont know how can you just give me the answer please.

OpenStudy (anonymous):

No sorry, thats not really the point of this site. Both sides can be divided by 6 to isolate the a^2. So now you have a^2=486/6.

OpenStudy (anonymous):

I dont know how to do that god.......Im to sleepy to remember / do it.

OpenStudy (anonymous):

It's done for you...

OpenStudy (anonymous):

Its 81

OpenStudy (anonymous):

Yep. Nice work.

OpenStudy (anonymous):

Thx.